OpenStudy (mackenzie2013):

What is the next term in the following sequence? 54, -18, 6, ... 3 2 -3 -2

OpenStudy (anonymous):

it would be -2.. you divide by -3

OpenStudy (anonymous):

r=-18/54=-1/3 in a G.P. \[tn=a r ^{n-1}\] a=54,n=4 calculate t4